OpenStudy (anonymous):

At a fair i bought 6 erasers and 2 pens for $15. If i could have bought 4 more erasers for $14 than pens for $9, the price of a pen was????

OpenStudy (anonymous):

@AriPotta @HELP!!!! @Austin_Rain @Aperez @asianfever @hba

OpenStudy (anonymous):

@alewis98 @BlackChickSwag @chance327 @DLS @evergirl

OpenStudy (aripotta):

i don't understand the last part of your question. is it saying 4 erasers cost $14??

OpenStudy (anonymous):

thats the problem its an aptitude question

OpenStudy (anonymous):

Is it saying 4 more erasers (plus the ones she bought) which would make it 10 erasers = $ 14 ???

OpenStudy (anonymous):

i am not able to understand the ques

OpenStudy (anonymous):

@damien @e.cociuba @firefightere9 @gjhfdfg @Hayatcakes @HB0315

OpenStudy (anonymous):

6e + 2p = 14 I am sorry...I can't get past this :(

OpenStudy (anonymous):

it's ok

OpenStudy (anonymous):

@Mimi_x3 @myko @n14r96

OpenStudy (aripotta):

the question's worded weird

OpenStudy (anonymous):

yes it is....confusing

OpenStudy (anonymous):

thank u ppl jus got the answer by my own

OpenStudy (aripotta):

cool. can you tell us the answer?

OpenStudy (aripotta):

just curious

OpenStudy (anonymous):

hmm.. yea

OpenStudy (anonymous):

the ques says that 6 erasers and 2 pens is for $15 so let us consider the cost of eraser and pen be E and P respectively => 6E+2P = 15 then , it says 4 more erasers for $14 than pens for $9 that is , if the number of pens is 'n' then the number of erasers is '(4+n)' nP=9 ,(4+n)E=14